# UniHash: Unifying Pointwise and Pairwise Hashing Paradigms for Seen and Unseen Category Retrieval

### v2

@@ Abstract
Effective retrieval across both seen and unseen categories is crucial for modern image retrieval systems. Retrieval on seen categories ensures precise recognition of known classes, while retrieval on unseen categories promotes generalization to novel classes with limited supervision. However, most existing deep hashing methods are confined to a single training paradigm, either pointwise or pairwise, where the former excels on seen categories and the latter generalizes better to unseen ones. To overcome this limitation, we propose Unified Hashing (UniHash), a dual-branch framework that unifies the strengths of both paradigms to achieve balanced retrieval performance across seen and unseen categories. UniHash consists of two complementary branches: a center-based branch following the pointwise paradigm and a pairwise branch following the pairwise paradigm. A novel hash code learning method is introduced to enable bidirectional knowledge transfer between branches, improving hash code discriminability and generalization. It employs a mutual learning loss to align hash representations and introduces a Split-Merge Mixture of Hash Experts (SM-MoH) module to enhance cross-branch exchange of hash representations. Theoretical analysis substantiates the effectiveness of UniHash, and extensive experiments on CIFAR-10, MSCOCO, and ImageNet demonstrate that UniHash consistently achieves state-of-the-art performance in both seen and unseen image retrieval scenarios.
@@ 1 Introduction
Hashing has become a prominent solution for large-scale image retrieval due to its efficiency in both computation and storage [20, 16, 2, 22, 18, 38]. It maps high-dimensional visual features into compact binary codes while preserving semantic similarity in the Hamming space [3, 42, 27, 15]. Recent deep hashing methods have achieved state-of-the-art performance by jointly optimizing feature extraction and hash code learning in an end-to-end fashion [43, 14, 34, 13, 26, 21, 5, 35, 31, 10]. Despite these advances, real-world retrieval systems often face queries from both seen categories encountered during training and unseen categories that emerge after deployment. Achieving balanced performance across these two settings is crucial for developing reliable retrieval systems, a challenge that existing hashing methods have largely overlooked.
@@
Deep supervised hashing methods can generally be divided into two paradigms: pointwise and pairwise. As shown in Fig. 1(a), recent pointwise methods ($e.g.,$ center-based methods)[43, 14, 34] introduce fixed or learnable hash centers that explicitly represent category prototypes. These methods establish representative hash centers for each category and train the network to align the hash codes of individual samples with their corresponding centers. Such pointwise supervision effectively enforces intra-class compactness [36, 41, 17, 33], yielding superior retrieval performance on seen categories. However, their strong reliance on predefined centers limits their generalization to unseen categories [30]. In contrast, as shown in Fig. 1(b), pairwise methods learn relative similarities among samples. Specifically, binary pairwise models [26, 21, 5] aim to pull similar pairs closer and push dissimilar pairs apart in the Hamming space. Triplet-based extensions [35, 8] introduce an anchor–positive–negative structure to capture relative ranking information, while listwise variants [23] further optimize global ranking metrics by jointly considering multiple items. This relational supervision flexibly preserves semantic consistency through feature similarity, thereby enhancing generalization to unseen categories.
As existing deep hashing methods are trained under a single supervision paradigm, either pointwise or pairwise, they struggle to achieve balanced image retrieval performance across seen and unseen categories [30]. To address this challenge, we propose Unified Hashing (UniHash), a dual-branch framework that unifies the strengths of both paradigms to achieve balanced image retrieval performance across seen and unseen categories. As illustrated in Fig. 1(c), UniHash consists of two complementary branches: a center-based branch corresponding to the pointwise paradigm and a pairwise branch corresponding to the pairwise paradigm. To enhance both the discriminability and generalization of hash codes, we introduce a novel hash code learning method that facilitates bidirectional knowledge exchange between two branches. Specifically, a mutual learning loss is employed to align hash code representations between two branches. Inspired by the Mixture of Experts (MoE) framework [29, 6, 7, 28], a Split-Merge Mixture of Hash Experts (SM-MoH) module is designed to promote cross-branch interaction in hash representation learning. Through this design, the center-based branch acquires relational cues from the pairwise branch, while the pairwise branch benefits from the compact semantic structure maintained by the center-based branch. Furthermore, we theoretically show that UniHash alleviates paradigm-specific discrepancies by enforcing the consistency between the two learning branches.
@@ 2 Related Works
Pointwise/Pairwise Training Paradigms. From the perspective of training paradigms, deep supervised hashing methods can be broadly categorized into pairwise and pointwise approaches. Early works such as Deep Hashing [25] and HashNet [5] follow the pairwise paradigm, which learns hash functions by modeling pairwise or triplet relationships. In contrast, pointwise methods such as CSQ [43], OrthoHash [14], and MDSH [34] define category specific hash centers, either pre-defined or learnable, to cluster semantically similar samples around a common center while pushing apart dissimilar ones. Recent studies indicate that pointwise methods perform better on seen categories, whereas pairwise methods generalize more effectively to unseen categories [30]. In this work, we aim to unify the strengths of both paradigms to achieve balanced retrieval performance across seen and unseen categories.
Deep Mutual Learning (DML). DML [44] is initially introduced as a collaborative training strategy in which multiple networks learn jointly by aligning their predicted class probabilities to improve generalization. Building on this idea, subsequent work extended DML to visual object tracking [46], where lightweight networks mutually supervise each other during offline training to enhance feature representation and tracking accuracy. Unlike prior approaches [44, 11, 45, 46] that train multiple networks under identical objectives, we enable two branches with distinct supervision paradigms to iteratively guide each other through hash code similarity for joint hash function optimization.
Mixture of Experts (MoE). The MoE framework [29] has recently been applied to large-scale neural networks, leveraging multiple specialized experts and a top-k gating mechanism for efficient computation. It has since been extended to multi-task learning in computer vision, such as AdaMV-MoE [6], which introduces task-specific routing and adaptive expert selection for diverse recognition tasks. Inspired by the effectiveness of MoE in modeling task-dependent information [29, 6, 7, 28], we propose the Split-Merge Mixture of Hash Experts (SM-MoH) module within our framework. Unlike standard MoE, SM-MoH adopts a split-merge routing mechanism: branch-specific gates independently select experts to capture distinct supervisory cues, while shared experts merge their outputs to ensure transformation consistency across branches. This design enhances cross-branch interaction in hash representation learning and improves the overall quality of hash codes.
@@ 3.1.1 Mutual Learning Loss
$P_{i,c}=\frac{\exp(\sqrt{q}\cos(u_{i}^{c},h_{c}))}{\sum_{m=1}^{C}\exp(\sqrt{q}\cos(u_{i}^{c},h_{m}))}.$ (3)
@@ 3.1.2 SM-MoH Module
Design Comparison. Unlike conventional MoE architectures that separate feature transformation and task-specific heads in Fig. 3(a), SM-MoH integrates expert transformation directly into the hashing process, as illustrated in Fig. 3(b)–(c). Each expert thus acts as an independent hashing pathway, producing semantically meaningful codes that reflect both branch-specific and shared characteristics. The final binary hash codes are computed as:
$b_{n}^{s}=\operatorname{sign}(u_{n}^{s}).$ (11)
This split–merge mechanism serves as an implicit bridge for semantic alignment, enhancing the discriminability of the resulting hash codes.
@@ 3.2.1 Preliminaries: Assumptions and Capacity Terms
Assumption 1 (Boundedness, Sparsity & Lipschitz).
@@
Assumption 2 (Center Separation).
@@
Assumption 3 (Semantic Manifolds).
The data reside on a union of semantic manifolds $\mathcal{M}=\bigcup_{t}\mathcal{M}_{t}$ with within-class diameter $\delta$ and inter-class separation $\Delta\gg\delta$ under a semantic metric. Under Assumption 2 and 3, quantization error decays exponentially in $q$:
@@
Assumption 4 (Mutual Consistency).
@@ 3.2.2 UniHash Superiority Analysis
Proposition 1 (Irreducible Structural Discrepancy).
@@
Theorem 1 (Discrepancy Elimination via UniHash).
@@
By Proposition 1, any single-branch model has a paradigm-induced error floor $E_{struct}>0$. The coupling loss $L_{M}$ in UniHash enforces branch agreement, replacing this floor by a consistency term $C_{cons}=\mathcal{O}(\tau/\sqrt{n})$. Together with the statistical and quantization terms, the UniHash risk satisfies
@@ 4 Experiments
Datasets. Following prior works [43, 14, 34, 26, 21, 5, 35, 31, 10], we evaluate performance on CIFAR-10 [19], ImageNet [9], and MSCOCO [24] for category-level retrieval. Evaluation metrics include mean average precision (mAP) and precision-recall curves. We report mAP@1000 for CIFAR-10 and ImageNet, and mAP@5000 for MSCOCO.
Training Setup. Following [43, 14, 34], we use a pre-trained ResNet-50 [12] as the backbone network $\phi(\cdot)$ and append a 4096-dimensional fully-connected ReLU layer [1] after the 2048-dimensional pooled feature to generate the base features for hashing. These features are processed by a Split-Merge Mixture of Hash Experts (SM-MoH) module, consisting of $m$ shared expert networks $\{E_{i}\}_{i=1}^{m}$, two gating networks $G_{\text{c}}$ and $G_{\text{p}}$, and hash centers $\{\mathbf{h}_{i}\}_{i=1}^{c}$ for the center stream. Input images are resized to $\times224$, and we use a mini-batch size of $N=64$. The model is trained for $T=100$ epochs to optimize the backbone $\phi$, experts $\{E_{i}\}$, and gating networks $G_{c}$, $G_{p}$, by jointly optimizing $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$. The final binary hash code is obtained as $\text{sign}(\mathbf{u}_{n}^{s})$, where $\mathbf{u}_{n}^{s}$ ($s\in\{{c},{p}\}$) are the refined continuous hash codes from the center and pairwise streams. Our model is implemented in PyTorch and trained on an NVIDIA RTX 4090 GPU using the RMSProp optimizer with a learning rate of 0.0001.
@@ 4.1 Closed-Set Retrieval Performance
Table 1: Comparison results of closed-set retrieval performance w.r.t. mAP on three datasets across different bit configurations. Bold values indicate the best performance, and underlined values indicate the second best performance.
@@
DSH[26] 0.7313 0.7402 0.7272 0.7179 0.7448 0.7585 0.7221 0.7573 0.7790
DPSH[21] 0.3098 0.3632 0.3638 0.6241 0.7626 0.7992 0.6239 0.6467 0.6322
HashNet[5] 0.8959 0.9115 0.8995 0.6024 0.7158 0.8071 0.7540 0.7331 0.7882
DTSH[35] 0.7783 0.7997 0.8312 0.6606 0.7803 0.8120 0.7702 0.8105 0.8233
GreedyHash[31] 0.3519 0.5350 0.6177 0.7394 0.7977 0.8243 0.7625 0.8033 0.8570
DPN[10] 0.7576 0.7901 0.8040 0.7987 0.8298 0.8394 0.7571 0.8227 0.8623
CSQ[43] 0.7861 0.7983 0.7989 0.8377 0.8750 0.8836 0.7509 0.8471 0.8610
OrthoHash[14] 0.9087 0.9297 0.9454 0.8540 0.8792 0.8936 0.7174 0.7675 0.8060
MDSH[34] 0.9455 0.9554 0.9607 0.8639 0.8863 0.9019 0.7542 0.8131 0.8143
CFBH[40] 0.7187 0.7397 0.7580 0.8120 0.8572 0.8950 0.6807 0.8325 0.8521
@@
Table 2: Retrieval under seen/unseen splits. mAP for Seen@Seen, Seen@All, Unseen@Unseen, and Unseen@All on 20%-unseen CIFAR-10 and 15%-unseen ImageNet/MSCOCO. Bold: best; Underlined: second best.
@@
CSQ [43] 95.2/95.1 95.1/95.0 95.0/94.9 85.3/85.0 89.1/88.7 90.6/90.2 65.0/79.7 68.8/81.9 70.1/82.4
OrthoHash [14] 94.6/92.8 95.1/93.2 95.7/93.7 88.2/87.7 90.3/89.8 91.5/90.9 62.7/78.7 65.8/80.5 68.5/82.4
MDSH [34] 92.8/92.3 92.5/92.0 92.8/89.9 75.5/75.3 78.8/78.5 81.9/81.1 64.2/80.1 70.3/81.2 70.9/83.3
CHN [4] 95.8/95.7 95.9/95.5 95.6/90.1 86.4/86.0 88.9/88.5 90.1/89.5
DSH [26] 92.6/91.9 92.5/91.8 92.9/92.0 68.8/67.9 79.5/78.2 85.3/84.0 46.8/64.6 51.1/69.7 51.4/71.8
HashNet [5] 80.5/80.1 92.7/92.2 94.3/93.4 49.1/48.9 75.8/75.0 86.5/85.6 52.4/74.8 59.9/79.9 64.6/82.4
Ours 97.3/97.0 97.2/97.1 97.2/97.0 89.9/88.3 92.2/89.9 92.4/91.5 70.3/81.0 72.5/82.3 73.2/83.9
@@
CSQ [43] 77.3/7.2 73.1/10.2 75.7/9.4 38.1/19.6 51.7/36.2 55.8/42.2 46.9/50.1 51.2/54.0 52.9/55.8
OrthoHash [14] 82.0/31.6 83.0/36.0 82.1/36.3 41.3/25.1 50.7/34.8 59.9/40.5 49.2/54.7 50.9/55.9 53.3/58.1
MDSH [34] 76.3/18.8 79.4/35.0 81.4/45.2 28.3/10.0 31.8/14.8 41.1/21.1 48.6/55.9 50.5/56.2 53.4/57.4
CHN [4] 73.2/6.6 73.2/11.6 78.0/39.0 37.0/18.0 46.9/30.8 56.2/39.0
DSH [26] 74.5/11.0 70.8/18.3 72.4/20.0 42.6/20.9 49.6/26.8 56.1/34.6 41.1/45.0 44.1/47.1 46.3/50.2
HashNet [5] 78.8/9.4 77.7/24.7 80.8/38.2 42.3/18.9 57.5/32.3 71.2/44.8 44.5/45.9 51.8/55.8 55.9/59.2
Ours 84.7/35.3 85.2/36.6 85.6/49.8 44.7/28.0 59.1/38.0 71.3/48.2 53.2/58.7 57.9/60.3 58.4/62.1
We compare our method with nine deep hashing algorithms: five pointwise methods (DPN [10], GreedyHash [31], CSQ [43], OrthoHash [14], MDSH [34]), three pairwise methods (DSH [26], DPSH [21], HashNet [5]), and one tripletwise method (DTSH [35]). Table 1 reports the Mean Average Precision (mAP) results for image retrieval. We adopt ResNet-50 as the backbone for all compared methods, including DSH, DPSH, DTSH, HashNet, GreedyHash, DPN, CSQ, OrthoHash, MDSH, and our proposed UniHash. Compared to these methods, our method achieves mAP improvements of 1.74%, 1.21%, and 0.87% on MSCOCO, CIFAR-10, and ImageNet, respectively, averaged across different code lengths.

### v3

@@ Abstract
Effective retrieval across both seen and unseen categories is crucial for modern image retrieval systems. Retrieval on seen categories ensures precise recognition of known classes, while retrieval on unseen categories promotes generalization to novel classes with limited supervision. However, most existing deep hashing methods are confined to a single training paradigm, either pointwise or pairwise, where the former excels on seen categories and the latter generalizes better to unseen ones. To overcome this limitation, we propose Unified Hashing (UniHash), a dual-branch framework that unifies the strengths of both paradigms to achieve balanced retrieval performance across seen and unseen categories. UniHash consists of two complementary branches: a center-based branch following the pointwise paradigm and a pairwise branch following the pairwise paradigm. A novel hash code learning method is introduced to enable bidirectional knowledge transfer between branches, improving hash code discriminability and generalization. It employs a mutual learning loss to align hash representations and introduces a Split-Merge Mixture of Hash Experts (SM-MoH) module to enhance cross-branch exchange of hash representations. Theoretical analysis substantiates the effectiveness of UniHash, and extensive experiments on CIFAR-10, MSCOCO, and ImageNet demonstrate that UniHash consistently achieves state-of-the-art performance in both seen and unseen image retrieval.
@@ 1 Introduction
Hashing has become a prominent solution for large-scale image retrieval due to its efficiency in both computation and storage (Li et al., 2011; Jiang et al., 2025; Cao et al., 2025; Liang et al., 2024; Kong et al., 2024; Wu et al., 2023). It maps high-dimensional visual features into compact binary codes while preserving semantic similarity in the Hamming space (Cao et al., 2018; Yang et al., 2018; Qin et al., 2018; Jégou et al., 2008). Recent deep hashing methods have achieved state-of-the-art performance by jointly optimizing feature extraction and hash code learning in an end-to-end fashion (Yuan et al., 2020; Hoe et al., 2021; Wang et al., 2023; He et al., 2024; Liu et al., 2016; Li et al., 2016; Cao et al., 2017b; Wang et al., 2017; Su et al., 2018; Fan et al., 2020). Despite these advances, real-world retrieval systems often face queries from both seen categories encountered during training and unseen categories that emerge after deployment. Achieving balanced performance across two settings is crucial for developing reliable retrieval systems, a challenge that existing hashing methods have overlooked.
@@
Deep supervised hashing methods can generally be divided into two paradigms: pointwise and pairwise. As shown in Fig. 1(a), recent pointwise methods ($e.g.,$ center-based methods)(Yuan et al., 2020; Hoe et al., 2021; Wang et al., 2023) introduce fixed or learnable hash centers that explicitly represent category prototypes. These methods establish representative hash centers for each category and train the network to align the hash codes of individual samples with their corresponding centers. Such pointwise supervision effectively enforces intra-class compactness (Wei et al., 2022; Xuan and Zhang, 2021; Kobayashi, 2021; Wang and Zhang, 2023), yielding superior retrieval performance on seen categories. However, their strong reliance on predefined centers limits their generalization to unseen categories (Shen et al., 2025). In contrast, as shown in Fig. 1(b), pairwise methods learn relative similarities among samples. Specifically, binary pairwise models (Liu et al., 2016; Li et al., 2016; Cao et al., 2017b) aim to pull similar pairs closer and push dissimilar pairs apart in the Hamming space. Triplet-based extensions (Wang et al., 2017; Deng et al., 2018) introduce an anchor–positive–negative structure to capture relative ranking information, while listwise variants (Liang et al., 2021) further optimize global ranking metrics by jointly considering multiple items. This relational supervision flexibly preserves semantic consistency through feature similarity, thereby enhancing generalization to unseen categories.
As existing deep hashing methods are trained under a single supervision paradigm, either pointwise or pairwise, they struggle to achieve balanced image retrieval performance across seen and unseen categories (Shen et al., 2025). To address this challenge, we propose Unified Hashing (UniHash), a dual-branch framework that unifies the strengths of both paradigms to achieve balanced image retrieval performance across seen and unseen categories. As illustrated in Fig. 1(c), UniHash consists of two complementary branches: a center-based branch corresponding to the pointwise paradigm and a pairwise branch corresponding to the pairwise paradigm. To enhance both the discriminability and generalization of hash codes, we introduce a novel hash code learning method that facilitates bidirectional knowledge exchange between two branches. Specifically, a mutual learning loss is employed to align hash code representations between two branches. Inspired by the Mixture of Experts (MoE) framework (Shazeer et al., 2017; Chen et al., 2023, 2022; Riquelme et al., 2021), a Split-Merge Mixture of Hash Experts (SM-MoH) module is designed to promote cross-branch interaction in hash representation learning. Through this design, the center-based branch acquires relational cues from the pairwise branch, while the pairwise branch benefits from the compact semantic structure maintained by the center-based branch. Furthermore, we theoretically show that UniHash alleviates paradigm-specific discrepancies by enforcing the consistency between the two learning branches.
@@ 2 Related Works
Pointwise/Pairwise Training Paradigms. From the perspective of training paradigms, deep supervised hashing methods can be broadly categorized into pairwise and pointwise approaches. Early works such as Deep Hashing (Liong et al., 2015) and HashNet (Cao et al., 2017b) follow the pairwise paradigm, which learns hash functions by modeling pairwise or triplet relationships. In contrast, pointwise methods such as CSQ (Yuan et al., 2020), OrthoHash (Hoe et al., 2021), and MDSH (Wang et al., 2023) define category specific hash centers, either pre-defined or learnable, to cluster semantically similar samples around a common center while pushing apart dissimilar ones. Recent studies indicate that pointwise methods perform better on seen categories, whereas pairwise methods generalize more effectively to unseen categories (Shen et al., 2025). In this work, we aim to unify the strengths of both paradigms to achieve balanced retrieval performance across seen and unseen categories.
Deep Mutual Learning (DML). DML (Zhang et al., 2018) is initially introduced as a collaborative training strategy in which multiple networks learn jointly by aligning their predicted class probabilities to improve generalization. Building on this idea, subsequent work extended DML to visual object tracking (Zhao et al., 2021), where lightweight networks mutually supervise each other during offline training to enhance feature representation and tracking accuracy. Unlike prior approaches (Zhang et al., 2018; Guo et al., 2022; Zhao and Han, 2021; Zhao et al., 2021) that train multiple networks under identical objectives, we enable two branches with distinct supervision paradigms to iteratively guide each other for joint hash function optimization.
Mixture of Experts (MoE). The MoE framework (Shazeer et al., 2017) has recently been applied to large-scale neural networks, leveraging multiple specialized experts and a top-k gating mechanism for efficient computation. It has since been extended to multi-task learning in computer vision, such as AdaMV-MoE (Chen et al., 2023), which introduces task-specific routing and adaptive expert selection for diverse recognition tasks. Inspired by the effectiveness of MoE in modeling task-dependent information (Shazeer et al., 2017; Chen et al., 2023, 2022; Riquelme et al., 2021), we propose the Split-Merge Mixture of Hash Experts (SM-MoH) module within our framework. Unlike standard MoE, SM-MoH adopts a split-merge routing mechanism: branch-specific gates independently select experts to capture distinct supervisory cues, while shared experts merge their outputs to ensure transformation consistency across branches. This design enhances cross-branch interaction in hash representation and improves the overall quality of hash codes.
@@ 3.1.1 Mutual Learning Loss
$P_{i,c}=\frac{\exp(\sqrt{q}\cos(u_{i}^{c},h_{c}))}{\sum_{r=1}^{C}\exp(\sqrt{q}\cos(u_{i}^{c},h_{r}))}.$ (3)
@@ 3.1.2 SM-MoH Module
Design Comparison. Unlike conventional MoE architectures that separate feature transformation and task-specific heads in Fig. 3(a), SM-MoH integrates expert transformation directly into the hashing process, as illustrated in Fig. 3(b)–(c). Each expert thus acts as an independent hashing pathway, producing semantically meaningful codes that reflect both branch-specific and shared characteristics. The final binary hash codes are computed as $b_{n}^{s}=\operatorname{sign}(u_{n}^{s})$.
We select the better-performing branch by mAP and use its binary codes for both queries and the database to perform Hamming-distance retrieval, where $s^{*}=\arg\max_{s\in\{c,p\}}\mathrm{mAP}(s)$ and $b^{s^{*}}=\mathrm{sign}\!\left(u^{s^{*}}\right)$. This split–merge mechanism serves as a bridge for semantic alignment, enhancing the discriminability of the resulting hash codes.
@@ 3.2.1 Preliminaries: Assumptions and Capacity Terms
Assumption 3.1 (Boundedness, Sparsity & Lipschitz).
@@
Assumption 3.2 (Center Separation).
@@
Assumption 3.3 (Semantic Manifolds).
The data reside on a union of semantic manifolds $\mathcal{M}=\bigcup_{t}\mathcal{M}_{t}$ with within-class diameter $\delta$ and inter-class separation $\Delta\gg\delta$ under a semantic metric. Under Assumption 3.2 and 3.3, quantization error decays exponentially in $q$:
@@
Assumption 3.4 (Mutual Consistency).
@@ 3.2.2 UniHash Superiority Analysis
Proposition 3.5 (Irreducible Structural Discrepancy).
@@
Theorem 3.6 (Discrepancy Elimination via UniHash).
@@
By Proposition 3.5, any single-branch model has a paradigm-induced error floor $E_{struct}>0$. The coupling loss $L_{M}$ in UniHash enforces branch agreement, replacing this floor by a consistency term $C_{cons}=\mathcal{O}(\tau/\sqrt{n})$. Together with the statistical and quantization terms, the UniHash risk satisfies
@@ 4 Experiments
Datasets. Following prior works (Yuan et al., 2020; Hoe et al., 2021; Wang et al., 2023; Liu et al., 2016; Li et al., 2016; Cao et al., 2017b; Wang et al., 2017; Su et al., 2018; Fan et al., 2020), we evaluate performance on CIFAR-10 (Krizhevsky and Hinton, 2009), ImageNet (Deng et al., 2009), and MSCOCO (Lin et al., 2014) for category-level retrieval. Evaluation metrics include mean average precision (mAP) and precision-recall curves. We report mAP@1000 for CIFAR-10 and ImageNet, and mAP@5000 for MSCOCO.
Training Setup. Following (Yuan et al., 2020; Hoe et al., 2021; Wang et al., 2023), we use a pre-trained ResNet-50 (He et al., 2016) as backbone $\phi(\cdot)$ and append a 4096-dimensional fully-connected ReLU layer (Agarap, 2019) after the 2048-dimensional pooled feature to generate the base features for hashing. These features are processed by a Split-Merge Mixture of Hash Experts (SM-MoH) module, consisting of $m$ shared expert networks $\{E_{i}\}_{i=1}^{m}$, two gating networks $G_{\text{c}}$ and $G_{\text{p}}$, and hash centers $\{\mathbf{h}_{i}\}_{i=1}^{c}$ for the center stream. Input images are resized to $\times224$, and we use a mini-batch size of $N=64$. The model is trained for $T=100$ epochs to optimize the backbone $\phi$, experts $\{E_{i}\}$, and gating networks $G_{c}$, $G_{p}$, by jointly optimizing $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$. The final binary hash code is obtained as $\text{sign}(\mathbf{u}_{n}^{s})$, where $\mathbf{u}_{n}^{s}$ ($s\in\{{c},{p}\}$) are refined continuous codes from center and pairwise streams. Our model is implemented in PyTorch and trained on NVIDIA RTX 4090 GPU using the RMSProp optimizer with learning rate of 0.0001.
@@ 4.1 Closed-Set Retrieval Performance
Table 1: Closed-set retrieval mAP on three datasets under different bit lengths. Bold indicates best performance; Underline indicates second best performance. Pairwise methods are shaded. DAHNet- denotes DAHNet without class-label supervision.
@@
DSH (Liu et al., 2016) 0.7313 0.7402 0.7272 0.7179 0.7448 0.7585 0.7221 0.7573 0.7790
DPSH (Li et al., 2016) 0.3098 0.3632 0.3638 0.6241 0.7626 0.7992 0.6239 0.6467 0.6322
HashNet (Cao et al., 2017b) 0.8959 0.9115 0.8995 0.6024 0.7158 0.8071 0.7540 0.7331 0.7882
DTSH (Wang et al., 2017) 0.7783 0.7997 0.8312 0.6606 0.7803 0.8120 0.7702 0.8105 0.8233
GreedyHash (Su et al., 2018) 0.3519 0.5350 0.6177 0.7394 0.7977 0.8243 0.7625 0.8033 0.8570
DPN (Fan et al., 2020) 0.7576 0.7901 0.8040 0.7987 0.8298 0.8394 0.7571 0.8227 0.8623
CSQ (Yuan et al., 2020) 0.7861 0.7983 0.7989 0.8377 0.8750 0.8836 0.7509 0.8471 0.8610
OrthoHash (Hoe et al., 2021) 0.9087 0.9297 0.9454 0.8540 0.8792 0.8936 0.7174 0.7675 0.8060
MDSH (Wang et al., 2023) 0.9455 0.9554 0.9607 0.8639 0.8863 0.9019 0.7542 0.8131 0.8143
AGMH (Lu et al., 2023) 0.7831 0.7925 0.8148 0.8535 0.8568 0.8471 0.7031 0.6283 0.6577
CFBH (Xiang et al., 2024) 0.7187 0.7397 0.7580 0.8120 0.8572 0.8950 0.6807 0.8325 0.8521
DAHNet- (Jiang et al., 2024) 0.8932 0.9087 0.9125 0.8325 0.8831 0.8921 0.7481 0.7993 0.8053
@@
Table 2: Retrieval under seen/unseen splits. mAP (%) for Seen@Seen, Seen@All, Unseen@Unseen, and Unseen@All on 20%-unseen CIFAR-10 and 15%-unseen ImageNet/MSCOCO. Bold: best; Underlined: second or third best. Pairwise methods are shaded. DAHNet denotes the label-supervised variant of DAHNet.
@@
CSQ (Yuan et al., 2020) 95.2/95.1 95.1/95.0 95.0/94.9 85.3/85.0 89.1/88.7 90.6/90.2 65.0/79.7 68.8/81.9 70.1/82.4
OrthoHash (Hoe et al., 2021) 94.6/92.8 95.1/93.2 95.7/93.7 88.2/87.7 90.3/89.8 91.5/90.9 62.7/78.7 65.8/80.5 68.5/82.4
MDSH (Wang et al., 2023) 92.8/92.3 92.5/92.0 92.8/89.9 75.5/75.3 78.8/78.5 81.9/81.1 64.2/80.1 70.3/81.2 70.9/83.3
CHN (Cao et al., 2017a) 95.8/95.7 95.9/95.5 95.6/90.1 86.4/86.0 88.9/88.5 90.1/89.5
DSH (Liu et al., 2016) 92.6/91.9 92.5/91.8 92.9/92.0 68.8/67.9 79.5/78.2 85.3/84.0 46.8/64.6 51.1/69.7 51.4/71.8
HashNet (Cao et al., 2017b) 80.5/80.1 92.7/92.2 94.3/93.4 49.1/48.9 75.8/75.0 86.5/85.6 52.4/74.8 59.9/79.9 64.6/82.4
ADSH (Jiang and Li, 2018) 95.3/67.0 95.0/69.2 93.8/61.7 83.5/82.7 89.7/88.6 90.4/89.3 75.1/66.7 79.7/71.3 80.0/68.3
AGMH (Lu et al., 2023) 81.0/41.5 79.4/12.5 81.0/5.5 85.8/85.2 87.4/86.0 84.0/82.6 69.0/58.2 59.7/38.8 64.1/39.7
DAHNet (Jiang et al., 2024) 94.8/67.5 93.7/63.2 92.2/58.6 86.1/85.3 88.8/87.5 88.6/87.5
Ours 97.3/97.0 97.2/97.1 97.2/97.0 89.9/88.3 92.2/89.9 92.4/91.5 79.3/81.0 81.5/82.3 85.2/83.9
@@
CSQ (Yuan et al., 2020) 77.3/7.2 73.1/10.2 75.7/9.4 38.1/19.6 51.7/36.2 55.8/42.2 46.9/50.1 51.2/54.0 52.9/55.8
OrthoHash (Hoe et al., 2021) 82.0/31.6 83.0/36.0 82.1/36.3 41.3/25.1 50.7/34.8 59.9/40.5 49.2/54.7 50.9/55.9 53.3/58.1
MDSH (Wang et al., 2023) 76.3/18.8 79.4/35.0 81.4/45.2 28.3/10.0 31.8/14.8 41.1/21.1 48.6/55.9 50.5/56.2 53.4/57.4
CHN (Cao et al., 2017a) 73.2/6.6 73.2/11.6 78.0/39.0 37.0/18.0 46.9/30.8 56.2/39.0
DSH (Liu et al., 2016) 74.5/11.0 70.8/18.3 72.4/20.0 42.6/20.9 49.6/26.8 56.1/34.6 41.1/45.0 44.1/47.1 46.3/50.2
HashNet (Cao et al., 2017b) 78.8/9.4 77.7/24.7 80.8/38.2 42.3/18.9 57.5/32.3 71.2/44.8 44.5/45.9 51.8/55.8 55.9/59.2
ADSH (Jiang and Li, 2018) 79.0/78.7 74.4/59.3 75.1/75.1 42.3/21.8 55.4/30.7 58.0/33.4 41.1/44.2 50.9/52.7 51.6/52.2
AGMH (Lu et al., 2023) 77.0/77.2 75.4/75.4 73.2/73.2 36.4/16.1 44.2/21.1 48.9/25.6 41.1/42.5 37.5/37.5 40.0/40.0
DAHNet (Jiang et al., 2024) 76.2/73.9 77.0/76.9 75.8/75.1 37.2/16.7 46.2/20.5 59.6/34.8
Ours 84.7/82.3 85.2/85.1 85.6/83.8 44.7/28.0 59.1/38.0 71.3/48.2 53.2/58.7 57.9/60.3 58.4/62.1
We compare our method with nine deep hashing algorithms: five pointwise methods (DPN (Fan et al., 2020), GreedyHash (Su et al., 2018), CSQ (Yuan et al., 2020), OrthoHash (Hoe et al., 2021), MDSH (Wang et al., 2023)), three pairwise methods (DSH (Liu et al., 2016), DPSH (Li et al., 2016), HashNet (Cao et al., 2017b)), AGMH (Lu et al., 2023), DAHNet- (Jiang et al., 2024), and one tripletwise method (DTSH (Wang et al., 2017)). Table 1 reports the Mean Average Precision (mAP) results for image retrieval. We adopt ResNet-50 as the backbone for all compared methods, including pairwise and pointwise baselines, and our proposed UniHash. Compared to these methods, our method achieves mAP improvements of 1.74%, 1.21%, and 0.87% on MSCOCO, CIFAR-10, and ImageNet, respectively, averaged across different code lengths.